OpenStudy (dls):

Disc is at rest on rough horizontal plane,mass=M.The disc is displaced slightly in horizontal direction.Friction is large enough that slipping ceases.Time period of oscillation?

OpenStudy (dls):

@Mashy @yrelhan4 why do i need to tag you people again n again :/ come on ur own!!!

OpenStudy (dls):



OpenStudy (yrelhan4):

ruk jaa. :/

OpenStudy (dls):

\[\LARGE \frac{qQ}{4 \pi \epsilon_o r^2} \times R = \frac{3}{2} MR^2 \times \frac{a}{R}\] is my torque equation about centre of mass..maybe :/

OpenStudy (dls):

@Mashy ?

OpenStudy (anonymous):

wow thats intense.. :O.. whats i Q? :O

OpenStudy (dls):

iQ?

OpenStudy (dls):

Q is a fixed charge outside and q is the charge on disc

OpenStudy (anonymous):

whats is q ?? in your equation?!

OpenStudy (dls):

^^

OpenStudy (anonymous):

wtf?!? there are charges involved as welL!? :O

OpenStudy (dls):

lol..how is it supposed to "oscillate"

OpenStudy (anonymous):

i thought due to gravity.. wait now that think of it.. thats not possible is it :P

OpenStudy (dls):